Some improvement on this report!
The camera for BSM_South has been repaired and is on its way back down to Australia. While Peter Nelson has been graciously loaning us his ST-10XME system, it will be good to be back on the air with the original camera. We're starting to get caught up on data processing.
BSM_Hamren is up and running at Ocean View on the Big Island. Michael Linnolt's brother is monitoring the system and is able to let it run most of the night. We still have some power issues (it is off the grid), but getting closer to being able to run all night.
BSM_HQ has finally gotten all of the bugs worked out of its new ACP V8 software. It is running very smoothly now on clear nights, thanks to the efforts of Helmar Adler.
BSM_Berry is still off the air until the mount is repaired. No time estimate at this point.
The replacement mount for BSM_NM has been installed by Bill Stein. He will be doing polar alignment this week, and we should be back into operation shortly thereafter.
coker30 is running nominally, just having poor weather.
OC61 is running nominally.
TMO61 is awaiting repair of the dome slit. We hope that will happen shortly.
SRO has had dome slit problems again. John Gross is out of town this week, but hopes to work on it this coming weekend.
So we have more systems operational now than on the last report, and a good chance that additional systems will be back in operation within a week or so.